OpenStudy (anonymous):

A manufacturer of light bulbs estimates that the fraction F(t) of bulbs that remain burning after t week is given by F(t)=e^(-kt) where k is a postive constant. Suppose twice as many bulbs are burning after 5 weeks as after 9 weeks. Find k and determine the fraction of bulbs still burning after 7 weeks. I can't figure out how to get k all I have so far is 2(e^(-5k))=e^(-9k) but I'm beginning to question whether that is right

OpenStudy (anonymous):

I did 2=e^(-9k)/e^(-5k) to get 2=e^(-4k) then took ln of both sides got ln2=-4k, idk if that is right

OpenStudy (anonymous):

im a tard shouldnt it be 2e^-9k=e^-5k ?

OpenStudy (anonymous):

yessss!

OpenStudy (anonymous):

then the answer makes sense now lol

OpenStudy (anonymous):

Lol Yess! it makes more sense xDDD :)

OpenStudy (anonymous):

did you get it or your still blah..Lol

OpenStudy (anonymous):

took my like a day to get this problem took you like 5 seconds the answer is ln2/4 right after that the rest of the problem is cake

OpenStudy (anonymous):

YESS!!!! wow.. a day?

OpenStudy (anonymous):

well I went to sleep on it picked up in the morning

OpenStudy (anonymous):

which was 5 minutes ago lol

OpenStudy (anonymous):

oh. :p